OpenStudy (anonymous):

If the volume of a sphere is 36π cubic units, what is the radius

OpenStudy (anonymous):

\[V=\frac{4}{3} \pi r^3\] Volume of a sphere We know that \[V=36 \pi\] so \[36 \pi=\frac{4}{3} \pi r^3\] Solve for r